# Information Limits of Joint Community Detection and Finite Group Synchronization

### v1

@@ Appendix B Proof of lemmas
$\mathbb{E}\left[t^{r}\right]\leq e^{-pN(1-t)}$
@@
$\mathbb{E}\left[\left(\frac{Ma}{b}\right)^{\alpha(r-r^{*})}\right]$ $=\mathbb{E}\left[\left(\frac{Ma}{b}\right)^{\alpha r}\right]\cdot\mathbb{E}\left[\left(\frac{Ma}{b}\right)^{-\alpha r^{*}}\right]$
$\leq\exp\left(-qN_{\Delta}\left(1-\left(\frac{Ma}{b}\right)^{\alpha}\right)\right)\cdot\exp\left(-pN_{\Delta}\left(1-\left(\frac{Ma}{b}\right)^{-\alpha}\right)\right)$
$=\exp\left(-\frac{N_{\Delta}\log n}{n}\left(a+b-a\left(\frac{Ma}{b}\right)^{-\alpha}-b\left(\frac{Ma}{b}\right)^{\alpha}\right)\right)$
$=\exp\left(-\left(\frac{N_{\Delta}\log n}{n}\right)f_{1}(\alpha)\right)$ (95)
@@
$f_{1}(\alpha):=a+b-a\left(\frac{Ma}{b}\right)^{-\alpha}-b\left(\frac{Ma}{b}\right)^{\alpha}.$
@@
$\mathbb{E}\left[\left(\frac{Ma}{b}\right)^{\alpha(r-r^{*})}\cdot M^{-(1-o(1))r}\right]$ $=\mathbb{E}\left[\left(\frac{Ma}{b}\right)^{\alpha r}\cdot M^{-(1-o(1))r}\right]\cdot\mathbb{E}\left[\left(\frac{Ma}{b}\right)^{-\alpha r^{*}}\right]$
$\leq\exp\left(-qN_{\Delta}\left(1-\left(\frac{Ma}{b}\right)^{\alpha}\cdot M^{-(1-o(1))}\right)\right)\cdot\exp\left(-pN_{\Delta}\left(1-\left(\frac{Ma}{b}\right)^{-\alpha}\right)\right)$
$=\exp\left(-\frac{N_{\Delta}\log n}{n}\left(a+b-a\left(\frac{Ma}{b}\right)^{-\alpha}-b\left(\frac{Ma}{b}\right)^{\alpha}\cdot M^{-(1-o(1))}\right)\right)$
$=\exp\left(-\left(\frac{N_{\Delta}\log n}{n}\right)f_{2}(\alpha)\right)$ (96)
@@
$f_{2}(\alpha):=a+b-a\left(\frac{Ma}{b}\right)^{-\alpha}-b\left(\frac{Ma}{b}\right)^{\alpha}\cdot M^{-(1-o(1))}.$
@@
$f_{2}(\alpha)\leq a+b-2\sqrt{\frac{ab}{M^{1-o(1)}}}=a+b-(2+o(1))\sqrt{\frac{ab}{M}}=O(1)$
@@
$a\left(\frac{Ma}{b}\right)^{-\alpha_{0}}=b\left(\frac{Ma}{b}\right)^{\alpha_{0}}\cdot M^{-(1-o(1))}\quad\Rightarrow\quad\left(\frac{Ma}{b}\right)^{\alpha_{0}}=\sqrt{\frac{aM^{1-o(1)}}{b}}.$
@@
$f_{1}(\alpha_{0})=-a-b+\sqrt{\frac{ab}{M^{1-o(1)}}}+\sqrt{abM^{1-o(1)}}=O(1).$
Now, plugging (95) and (96) into (94) and letting $\alpha=\alpha_{0}$ yields
@@
$=n^{-\Theta(\log n)}\cdot\exp\left(-\left(\frac{N_{\Delta}\log n}{n}\right)\cdot O(1)\right)+M^{2\Delta}\cdot\exp\left(-\left(\frac{N_{\Delta}\log n}{n}\right)\cdot f_{2}(\alpha_{0})\right)$
@@
$=M^{2\Delta}\cdot\exp\left(-\frac{N_{\Delta}\log n}{n}\left(a+b-(2+o(1))\sqrt{\frac{ab}{M}}\right)\right)$
@@ B.3 Proof of Lemma 56
$r_{1}^{\prime}\sim\textrm{Binom}\left(\frac{n}{2},\;p\right)$
@@
$\geq\sum_{r_{2}=\delta}^{n/2}\mathcal{P}(r_{2})\cdot\mathcal{P}\left(r_{1}^{\prime}=r_{2}-\delta\right)\cdot M^{-r_{2}}$ (97)
@@
$k^{*}=\operatorname*{arg\,max}\limits_{0\leq k\leq\frac{n}{2}-\delta}f(k)=\Theta(\log n)$
@@
$=\frac{\left(\frac{1}{2}-\frac{k}{n}\right)\left(\frac{1}{2}-\frac{k-\delta}{n}\right)}{(k+1)(k+\delta+1)}\cdot\left(\frac{ab\cdot\log^{2}n}{M}\right)(1+o(1)).$ (101)
@@
$\frac{f(k+1)}{f(k)}\leq\left(\frac{\log^{2}n}{k^{2}}\right)\cdot\left(\frac{ab}{4M}\right)\cdot(1+o(1))=o(1)$
@@
$\frac{f(k+1)}{f(k)}\geq\frac{\log^{2}n}{(k+\delta+1)^{2}}\cdot\left(\frac{ab}{4M}\right)\cdot(1+o(1))=\omega(1)$
@@
Next, we consider each term separately. First, by using the binomial coefficient bound [31, Lemma 23.1] that $\binom{n}{k}\geq\frac{n^{k}}{k!}\left(1-\frac{k(k-1)}{2n}\right)$ , we have
@@
$=k\left[\log\left(\frac{n}{2}\right)-\log\left(\frac{k}{e}\right)-o(1)\right]$ (103)
@@
$=h(k)+\left(-\frac{a+b}{2}+1\right)\log n-o(\log n)$ (107)
@@ B.4 Proof of Lemma 59
$\geq\left(-\frac{a+b}{2}+\sqrt{\frac{ab}{M}}\right)\log n-o(\log n)$

### v2

@@ Appendix B Proof of lemmas
$\mathbb{E}\left[t^{r}\right]\leq e^{-pN(1-t)},$
@@
$\mathbb{E}\left[\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha(r-r^{*})}\right]$ $=\mathbb{E}\left[\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha r}\right]\cdot\mathbb{E}\left[\left(\frac{Ma(1-q)}{b(1-p)}\right)^{-\alpha r^{*}}\right]$
$\leq\exp\left(-qN_{\Delta}\left(1-\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha}\right)\right)\cdot\exp\left(-pN_{\Delta}\left(1-\left(\frac{Ma(1-q)}{b(1-p)}\right)^{-\alpha}\right)\right)$
$=\exp\left(-\frac{N_{\Delta}\log n}{n}\left(a+b-a\left(\frac{Ma(1-q)}{b(1-p)}\right)^{-\alpha}-b\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha}\right)\right)$
$=\exp\left(-\left(\frac{N_{\Delta}\log n}{n}\right)f_{1}(\alpha)\right),$ (104)
@@
$f_{1}(\alpha):=a+b-a\left(\frac{Ma(1-q)}{b(1-p)}\right)^{-\alpha}-b\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha}.$
@@
$\mathbb{E}\left[\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha(r-r^{*})}\cdot M^{-(1-o(1))r}\right]$
$=\mathbb{E}\left[\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha r}\cdot M^{-(1-o(1))r}\right]\cdot\mathbb{E}\left[\left(\frac{Ma(1-q)}{b(1-p)}\right)^{-\alpha r^{*}}\right]$
$\leq\exp\left(-qN_{\Delta}\left(1-\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha}\cdot M^{-(1-o(1))}\right)\right)\cdot\exp\left(-pN_{\Delta}\left(1-\left(\frac{Ma(1-q)}{b(1-p)}\right)^{-\alpha}\right)\right)$
$=\exp\left(-\frac{N_{\Delta}\log n}{n}\left(a+b-a\left(\frac{Ma(1-q)}{b(1-p)}\right)^{-\alpha}-b\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha}\cdot M^{-(1-o(1))}\right)\right)$
$=\exp\left(-\left(\frac{N_{\Delta}\log n}{n}\right)f_{2}(\alpha)\right),$ (105)
@@
$f_{2}(\alpha):=a+b-a\left(\frac{Ma(1-q)}{b(1-p)}\right)^{-\alpha}-b\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha}\cdot M^{-(1-o(1))}.$
@@
$f_{2}(\alpha)\leq a+b-2\sqrt{\frac{ab}{M^{1-o(1)}}}=a+b-(2+o(1))\sqrt{\frac{ab}{M}}=\Theta(1),$ (106)
@@
$a\left(\frac{Ma(1-q)}{b(1-p)}\right)^{-\alpha_{0}}=b\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha_{0}}\cdot M^{-(1-o(1))}\quad\Rightarrow\quad\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha_{0}}=\sqrt{\frac{aM^{1-o(1)}}{b}}.$
@@
$f_{1}(\alpha_{0})=-a-b+\sqrt{\frac{ab}{M^{1-o(1)}}}+\sqrt{abM^{1-o(1)}}=\Theta(1).$ (107)
Now, given (107) and (106), by letting $\alpha=\alpha_{0}$ we can upper bound (103) as
@@
$=n^{-\Theta(\log n)}\cdot\exp\left(-\left(\frac{N_{\Delta}\log n}{n}\right)\cdot\Theta(1)\right)+M^{2\Delta}\cdot\exp\left(-\left(\frac{N_{\Delta}\log n}{n}\right)f_{2}(\alpha_{0})\right)$
@@
$\leq M^{2\Delta}\cdot\exp\left(-\frac{N_{\Delta}\log n}{n}\left(a+b-(2+o(1))\sqrt{\frac{ab}{M}}\right)\right),$
@@ B.3 Proof of Lemma 65
$r_{1}^{\prime}\sim\textrm{Binom}\left(\frac{n}{2},\;p\right),$
@@
$\geq\sum_{r_{2}=\delta}^{n/2}\mathcal{P}(r_{2})\cdot\mathcal{P}\left(r_{1}^{\prime}=r_{2}-\delta\right)\cdot M^{-r_{2}},$ (108)
@@
$k^{*}=\operatorname*{arg\,max}\limits_{0\leq k\leq\frac{n}{2}-\delta}f(k)=\Theta(\log n).$
@@
$=\frac{\left(\frac{1}{2}-\frac{k}{n}\right)\left(\frac{1}{2}-\frac{k-\delta}{n}\right)}{(k+1)(k+\delta+1)}\cdot\left(\frac{ab\cdot\log^{2}n}{M}\right)(1+o(1)),$ (112)
@@
$\frac{f(k+1)}{f(k)}\leq\left(\frac{\log^{2}n}{k^{2}}\right)\cdot\left(\frac{ab}{4M}\right)\cdot(1+o(1))=o(1),$
@@
$\frac{f(k+1)}{f(k)}\geq\frac{\log^{2}n}{(k+\delta+1)^{2}}\cdot\left(\frac{ab}{4M}\right)\cdot(1+o(1))=\omega(1),$
@@
Next, we consider each term separately. First, by using the binomial coefficient bound [32, Lemma 23.1] that $\binom{n}{k}\geq\frac{n^{k}}{k!}\left(1-\frac{k(k-1)}{2n}\right)$, we have
@@
$=k\left[\log\left(\frac{n}{2}\right)-\log\left(\frac{k}{e}\right)-o(1)\right],$ (114)
@@
$=h(k)+\left(-\frac{a+b}{2}+1\right)\log n-o(\log n).$ (118)
@@ B.4 Proof of Lemma 68
$\geq\left(-\frac{a+b}{2}+\sqrt{\frac{ab}{M}}\right)\log n-o(\log n),$